# Interpreting Graph Inference with Skyline Explanations

### v2

@@ IV-B Diversified Skyline Explanations
${\sf CD}(G_{s},G_{s^{\prime}})=1-\frac{\mathbf{x}_{G_{s}}\cdot\mathbf{x}_{G_{s^{\prime}}}}{||\mathbf{x}_{G_{s}}||_{2}\cdot||\mathbf{x}_{G_{s^{\prime}}}||_{2}}$ (7)
@@
Diversification Algorithm. We next outline our diversified evaluation algorithm, denoted as $\mathsf{DSX}$ (pseudo-code shown in Appendix). It follows $\mathsf{ASX}$-$\mathsf{OP}$ and adopts onion peeling strategy to generate subgraphs in a stream. The difference is that when computing the $k$-skyline explanation, it includes a new replacement strategy when the marginal gain for the diversification function ${\sf DivS}(\cdot)$ is at least a factor of the score over the current solution ${\sf DivS}({\mathcal{G}}_{\epsilon})$. $\mathsf{DSX}$ terminates when a first $k$-skyline explanation is found.
@@
We show the above result by reducing the diversified evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the Streaming Submodular Maximization problem [47]. The problem maintains a size-$k$ set that optimizes a submodular function over a stream of data objects. We show that $\mathsf{DSX}$ adopts a consistent increment policy that ensures a ($\frac{1}{2}-\epsilon$)-approximation as in [47]. We present the detailed proof in Appendix.
@@ V Parallel Skyline Explanatory Querying
Parallel Algorithm. The parallel algorithm, denoted as $\mathsf{ParaSX}$, works with $m$ threads to evaluate a query set $Q$ with $n$ ${\mathsf{SXQ}}$ queries, involving a test node set $V_{T}$ (pseudocode given in Appendix). (1) Inter-Batch parallelization: It first partitions $Q$ into $N$ small batches to be evaluated in parallel, guided by a clustering strategy with a goal that for any pair of target nodes $v_{t}$ and $v_{t}^{\prime}$ from two different batches, their $L$-hop neighbors have minimized overlap (see “Workload Clustering”). (2) Batch Assignment. $\mathsf{ParaSX}$ then solves a makespan minimization problem to assign the $N$ batches to $m$ threads, with a dynamic scheduling strategy that prioritizes the processing of queries with “influencing” output nodes that have highest common $L$-hop neighbors for those in the same thread. This is to maximize the shared computation for edge peeling and verification. (3) Inter-query Pipelining: Each local thread then invokes $\mathsf{ASX}$-$\mathsf{OP}$ ($v_{t}$) to generate local $k$-explanations for each assigned node $v_{t}$. All the processing follows its local ranks of highly influential nodes, and automatically skips processing a node, if its information has been updated via peeling operations earlier by previous nodes. This processing strategy is enabled by maintaining a shared data structure that records nodes and their $L$-hop neighboring edges within each thread. The skyline computation can be further parallelized by parallel skyline query methods [48, 49].
(a) IPF Overall Score
(b) IGD Overall Score
(c) MS on $\mathsf{Cora}$
(d) MS on $\mathsf{PubMed}$
[…]
Parallelization Strategies. $\mathsf{ParaSX}$ adopts the two strategies below. (1) Workload Clustering. The workload partitioning is guided by solving an $N$-clustering $\Psi$ of $Q$ that maximizes the accumulated intra-cluster Jaccard similarity (J) [50] of the $L$-hop neighbors of the targeted nodes in each cluster. This induces query workloads in batches with targeted nodes sharing the most common $L$-hop neighbors. The accumulated pairwise intra-similarity of $\mathsf{SXQ}$ queries is defined as $\sum_{{\mathsf{SXQ}}^{k}_{1},{\mathsf{SXQ}}^{k}_{2}\in\Psi(t)}J(N^{L}(v_{t}^{1}),N^{L}(v_{t}^{2}))$, where $N^{L}(\cdot)$ denotes the $L$-hop neighbor set of a given node, and $v_{t}^{1}$ and $v_{t}^{2}$ denote the test nodes of ${\mathsf{SXQ}}^{k}_{1}$ and ${\mathsf{SXQ}}^{k}_{2}$, respectively. (2) Influential Node Prioritization. The Inter-query pipelining exploits the shared global data structure to dynamically update an influencing score of the nodes. Specifically, each thread begins by applying the $k$-core decomposition algorithm [51] to assign a core number to each node, reflecting the highest $j$ such that the node belongs to the $j$-core. Test nodes are then ranked in descending order of core number, prioritizing more “central” nodes in the thread. This ordering ensures that edge information from structurally important nodes is shared early, maximizing reusability (see more details in Appendix).
@@ VI-A Experimental Setup
Datasets. Used datasets are summarized in Table III. (1) $\mathsf{Cora}$ [52] and $\mathsf{PubMed}$ [53] are citation networks with a set of papers (nodes) and their citation relations (edges). Each node has a feature vector encoding the presence of a keyword from a dictionary. For both, we consider a node classification task that assigns a paper category to each node. (2) In $\mathsf{FacebookPage}$ [54], the nodes represent verified Facebook pages, and edges are mutual “likes”. The node features are extracted from the site descriptions. The task is multi-class classification, which assigns multiple site categories (politicians, governmental organizations, television shows, and companies) to a page. (3) $\mathsf{AmazonComputer}$ [55] is a product network. The nodes represent “Computer” products and an edge between two products encodes that the two products are co-purchased by the same customer. The node features are product reviews as bag-of-words. The task is to classify the product categories. (4) $\mathsf{OGBN\_arxiv}$ [56] is a citation network of Computer Science papers. Each paper comes with a 128-dimensional feature vector obtained by word embeddings from its title and abstract. The task is to classify the subject areas. (5) $\mathsf{BA\_Shapes}$ [6] is a billion-scale synthetic graph, augmented with house-shaped motifs. Each node belongs to one of four classes: top, middle, or bottom (of a house motif), or non-house (not part of any motif).
@@
[…]
@@
[…]
Evaluation Metrics. For all datasets and $\mathsf{GNNs}$, we select three common explainability measures ($\Phi$): ${\mathsf{fdl}}^{+}$, ${\mathsf{fdl}}^{-}$, and $\mathsf{conc}$. As most $\mathsf{GNN}$ explainers are not designed to generate explanations for multiple explainability measures, for a fair comparison, we employ three quality indicators (QIs)  [58, 59, 60, 61]. These quality indicators are widely-used to measure how good each result is in a multi-objective manner. Consider a set of $n$ $\mathsf{GNN}$ explainers, where each explainer $A$ reports a set of explanatory subgraphs ${\mathcal{G}}_{A}$.
(1) QI-1: Integrated Preference Function (IPF) [62]. IPF score unifies and compares the quality of non-dominated set solutions with a weighted linear sum function. We define a normalized IPF of an explanation ${\mathcal{G}}_{A}$ from each $\mathsf{GNN}$ explainer $A$ with a normalized single-objective score:
@@
(2) QI-2: Inverted Generational Distance (IGD) [63, 58], a most commonly used distance-based QI. It measures the distance from each solution to a reference set that contains top data points with theoretically achievable “ideal” values. We introduce IGD for explanations as follows. (a) We define a universal space ${\mathcal{G}}$ = $\bigcup_{i\in[1,n]}{\mathcal{G}}_{i}$ from all the participating $\mathsf{GNN}$ explainers, and for each explainability measure $\phi\in\Phi$, induces a reference set ${\mathcal{G}}_{\phi}^{k}\in{\mathcal{G}}$ with explanatory subgraphs having the top-$k$ values in $\phi$. (b) The normalized IGD of an explanation ${\mathcal{G}}_{A}$ from an explainer $A$ is defined as:
@@
(3) QI-3: Maximum Spread (MS) [58]. $\mathsf{MS}$ is a widely-adopted spread indicator that quantifies the range of the minimum and maximum values a solution can achieve in each objective. For a fair comparison, we introduce a normalized $\mathsf{MS}$ score using reference sets in QI-2. For each measure $\phi\in\Phi$, and an explanation ${\mathcal{G}}_{A}$, its normalized MS score on $\phi$ is computed as:
@@ VI-B Experimental Results
Exp-1: Overall Explainability. We evaluate the overall performance of the $\mathsf{GNN}$ explainers using QIs.
[…]
@@
QI-3: nMX Scores. Figures 3(c) and  3(d) visualize the normalized MS scores of the $\mathsf{GNN}$ explainers, over $\mathsf{Cora}$ and $\mathsf{PubMed}$, respectively, where $k$ = $10$. (1) $\mathsf{ASX}$-$\mathsf{OP}$ reports a large range and contributes to explanations that are either close or the exact optimal explanation, for each of the three measures. $\mathsf{ASX}$-$\mathsf{I}$ and $\mathsf{DSX}$ have comparable performance, yet with larger ranges due to its properly diversified solution. (2) $\mathsf{DSX}$, $\mathsf{ASX}$-$\mathsf{OP}$, and $\mathsf{ASX}$-$\mathsf{I}$ contribute to the optimal ${\mathsf{fdl}}^{+}$, ${\mathsf{fdl}}^{-}$, and ${\mathsf{conc}}$, respectively. On the other hand, each individual explainer performs worse for all measures, with a large gap. For example, in $\mathsf{Cora}$, $\mathsf{MOExp}$ only achieves up to 3% of the best explanation ($\mathsf{DSX}$) over ${\mathsf{fdl}}^{-}$.
@@
[…]
@@
(b) Visualization of Skyline Explanation ($\mathsf{AmazonComputer}$)
@@
(d) Visualization of Skyline Explanation ($\mathsf{FacebookPage}$)
[…]
Exp-3: Scalability. We report the impact of critical factors (i.e., number of explanatory subgraphs $k$, $\mathsf{GNN}$ classes, and number of $\mathsf{GNN}$ layers $L$) on the scalability of skyline explanation generation, using $\mathsf{Cora}$ dataset. Additional results about effectiveness by varying factors are in Appendix.
@@
Varying $\mathsf{GNN}$ classes and $L$. (1) Fixing $k$=$10$, we report the time cost of $3$-layer $\mathsf{GNN}$ explainers for $\mathsf{GCN}$, $\mathsf{GAT}$, and $\mathsf{GIN}$ over $\mathsf{Cora}$. As shown in Figure 4(c), all our skyline methods take the least time to explain $\mathsf{GNN}$-based classification. This is consistent with our observation that the verification of $\mathsf{GNN}$ is the most efficient among all classes, indicating an overall small verification cost. (2) We fix $k$=$10$ and report the time cost of $\mathsf{GCN}$ explainers, with the number of layers $L$ varied from 1 to 3 over $\mathsf{Cora}$. Figure 4(d) shows us that all our methods significantly outperform the competitors. The learning overheads of competitors remain their major bottleneck, while our algorithms, as post-hoc explainers without learning overhead, are more efficient. As expected, all our methods take a longer time to generate explanations for larger $L$, as more subgraphs need to be verified from larger induced $L$-hop neighbors.
@@
Diversified Explanation. A user is interested in finding “Why” a product $v_{1}$ is labeled “PC Gaming” by a $\mathsf{GCN}$. $\mathsf{GNN}$ explainers that optimize a single measure (e.g.,$\mathsf{GExp}$) return explanations (see $G_{\zeta}^{8}$ in Figure 5(a)) that simply reveals a fact that $v_{1}$ is co-purchased with “Components”, a “one-sided” interpretation. In contrast, $\mathsf{DSX}$ identifies an explanation that reveals more comprehensive interpretation with two explanatory subgraphs (both factual) $\{G_{\zeta}^{9},G_{\zeta}^{10}\}$ (Figure 5(a)), which reveal two co-purchasing patterns bridging $v_{1}$ with not only “Components” for “Monitors” (by $G_{\zeta}^{9}$), but also to “Accessories” of “Laptops”. Indeed, the former indicates gamers who prefer gaming PC with high-refresh rate monitors; and the latter indicates that $v_{1}$ is a gaming laptop in need of frequent maintenance with laptop accessories. This showcase that $\mathsf{DSX}$ provides more comprehensive explanation.
@@
Skyline vs. Multi-objective Explanation (Linear Combination). Our second case compares $\mathsf{DSX}$ and $\mathsf{PGExp}$ (shown in Figure 5(c)). The latter generates explanations by optimizing a single objective that combines two explanation measures (factuality and conciseness). We observe that $\mathsf{PGExp}$ generates small factual subgraphs, such as $G_{\zeta}^{11}$ for a test node $v_{2}$, yet relatively less informative. $\mathsf{DSX}$ generates a skyline explanation $\{G_{\zeta}^{12},G_{\zeta}^{13}\}$ that is able to interpret $v_{2}$ as “Politicians” not only due to its connection with “Governments” entities but also highly related to “TV shows” and “Companies”. Meanwhile, Figure 5(d) verifies that our skyline explanation covers most of the interpretable space with diverse and comprehensive explanations, while the SOTA explainers cluster their solutions in smaller regions.

### v3

@@ IV-B Diversified Skyline Explanations
[…]
@@
Diversification Algorithm. We next outline our diversified evaluation algorithm, denoted as $\mathsf{DSX}$ (pseudo-code shown in [47]). It follows $\mathsf{ASX}$-$\mathsf{OP}$ and adopts onion peeling strategy to generate subgraphs in a stream. The difference is that when computing the $k$-skyline explanation, it includes a new replacement strategy when the marginal gain for the diversification function ${\sf DivS}(\cdot)$ is at least a factor of the score over the current solution ${\sf DivS}({\mathcal{G}}_{\epsilon})$. $\mathsf{DSX}$ terminates when a first $k$-skyline explanation is found.
@@
We show the above result by reducing the diversified evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the Streaming Submodular Maximization problem [53]. The problem maintains a size-$k$ set that optimizes a submodular function over a stream of data objects. We show that $\mathsf{DSX}$ adopts a consistent increment policy that ensures a ($\frac{1}{2}-\epsilon$)-approximation as in [53]. We present the detailed proof in [47].
@@ V Parallel Skyline Explanatory Querying
Parallel Algorithm. The parallel algorithm, denoted as $\mathsf{ParaSX}$, works with $m$ threads to evaluate a query set $Q$ with $n$ $\mathsf{SXQ}$} queries, involving a test node set $V_{T}$ (pseudocode given in the full version [47]). (1) Inter-Batch parallelization: It first partitions $Q$ into $N$ small batches to be evaluated in parallel, guided by a clustering strategy with a goal that for any pair of target nodes $v_{t}$ and $v_{t}^{\prime}$ from two different batches, their $L$-hop neighbors have minimized overlap (see “Workload Clustering”). (2) Batch Assignment. $\mathsf{ParaSX}$ then solves a makespan minimization problem to assign the $N$ batches to $m$ threads, with a dynamic scheduling strategy that prioritizes the processing of queries with “influencing” output nodes that have highest common $L$-hop neighbors for those in the same thread. This is to maximize the shared computation for edge peeling and verification. (3) Inter-query Pipelining: Each local thread then invokes $\mathsf{ASX}$-$\mathsf{OP}$ ($v_{t}$) to generate local $k$-explanations for each assigned node $v_{t}$. All the processing follows its local ranks of highly influential nodes, and automatically skips processing a node, if its information has been updated via peeling operations earlier by previous nodes. This processing strategy is enabled by maintaining a shared data structure that records nodes and their $L$-hop neighboring edges within each thread. The skyline computation can be further parallelized by parallel skyline query methods [54, 55].
(a) IPF score
(b) IGD score
(c) MS score ($\mathsf{Cora}$)
(d) IPF score over subsets of $\Phi$ ($\mathsf{Tree\_Cycles}$)
Figure 3: Effectiveness of our skyline explanation vs. baseline explainers ($\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, $\mathsf{SAME}$).
Parallelization Strategies. $\mathsf{ParaSX}$ employs two strategies. (1) Workload Clustering. We partition query workloads by solving an $N$-clustering of $Q$ that maximizes intra-cluster Jaccard similarity of $L$-hop neighbors. This groups test nodes with overlapping neighborhoods, enabling efficient batched processing. (2) Influential Node Prioritization. We apply $k$-core decomposition to rank nodes by structural centrality. Test nodes with higher core numbers are processed first, ensuring that edge information from influential nodes is shared early across threads for maximal reuse.
@@ VI-A Experimental Setup
Datasets. Table III summarizes the datasets. (1) $\mathsf{Cora}$ [56] and $\mathsf{PubMed}$ [57] are citation networks where nodes (papers) are classified by category using keyword-based features. (2) $\mathsf{FacebookPage}$ ($\mathsf{FB}$[58] contains verified Facebook pages (nodes) linked by mutual likes, with features from page descriptions; the task is multi-class page classification. (3) $\mathsf{AmazonComputer}$ ($\mathsf{AMZ}$[59] is a co-purchase network of computer products with review-based features; the task is product category classification. (4) $\mathsf{OGBN\_arxiv}$ ($\mathsf{arxiv}$[60] is a citation network of CS papers with 128-d word-embedding features, classified by subject area. (5) $\mathsf{BA\_Shapes}$ ($\mathsf{BA}$[6] is a synthetic BA graph with house motifs; house nodes serve as human-crafted ground-truth explanations. (6) $\mathsf{Tree\_Cycles}$ ($\mathsf{Tree}$[6] is a synthetic tree graph with cycle motifs; cycle nodes serve as human-crafted ground-truth explanation [5].
@@
We compare against several state-of-the-art $\mathsf{GNN}$ explainers. (1) GNNExplainer ($\mathsf{GExp}$) learns edge and feature masks by maximizing mutual information between predictions on the original and masked graphs [6]. (2) PGExplainer ($\mathsf{PGExp}$) trains an MLP as a mask generator on node embeddings, also optimizing mutual information [10]. (3) $\mathsf{CF\textsuperscript{2}}$ learns feature and edge masks by optimizing a weighted combination of factual and counterfactual measures [8]. (4) $\mathsf{MOExp}$ produces Pareto-optimal explanations balancing simulatability and counterfactual relevance [11]; without constraint $k$, it may return up to $\sim$200 subgraphs. (5) SAME develops structure-aware Shapley-based explanations via connected substructures [34].
@@
Figure 4: Efficiency and scalability of our skyline explanation vs. SOTA explainers ($\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, $\mathsf{SAME}$).
Evaluation Metrics. For all datasets and $\mathsf{GNNs}$, we select three common explainability measures ($\Phi$): $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$. In addition, we adopt the explanation accuracy [6] ($\mathsf{acc}$) for the synthetic $\mathsf{Tree\_Cycles}$ dataset. $\mathsf{acc}$ measures the proportion of human-crafted ground-truth nodes present in the explanatory subgraph. For computational simplicity, we transformed each minimization measure ($\mathsf{conc}$}, $\mathsf{fdl}$$^{-}$) into its inverse counterpart, so that all measures are expressed as maximization objectives. As most $\mathsf{GNN}$ explainers are not designed to generate explanations for multiple explainability measures, for a fair comparison, we employ three quality indicators (QIs)  [62, 63, 64, 65]. These quality indicators are widely-used to measure how good each result is in a multi-objective manner. Consider a set of $n$ $\mathsf{GNN}$ explainers, where each explainer $A$ reports a set of explanatory subgraphs ${\mathcal{G}}_{A}$.
(1) QI-1: Integrated Preference Function (IPF) (Higher is better) [66]. IPF score unifies and compares the quality of non-dominated set solutions with a weighted linear sum function. We define a normalized IPF of an explanation ${\mathcal{G}}_{A}$ from each $\mathsf{GNN}$ explainer $A$ with a normalized single-objective score:
@@
(2) QI-2: Inverted Generational Distance (IGD) (Lower is better)  [67, 62], a most commonly used distance-based QI. It measures the distance from each solution to a reference set that contains top data points with theoretically achievable “ideal” values. We introduce IGD for explanations as follows. (a) We define a universal space $\mathcal{G}$} = $\bigcup_{i\in[1,n]}{\mathcal{G}}_{i}$ from all the participating $\mathsf{GNN}$ explainers, and for each $\phi\in\Phi$, induces a reference set ${\mathcal{G}}_{\phi}^{k}\in{\mathcal{G}}$ with explanatory subgraphs having the top-$k$ values in $\phi$. (b) The normalized IGD of an explanation ${\mathcal{G}}_{A}$ from an explainer $A$ is defined as:
@@
(3) QI-3: Maximum Spread (MS) (Higher is better) [62]. $\mathsf{MS}$ is a widely-adopted spread indicator that quantifies the range of the minimum and maximum values a solution can achieve in each objective. For a fair comparison, we introduce a normalized $\mathsf{MS}$ score using reference sets in QI-2. For each measure $\phi\in\Phi$, and an explanation ${\mathcal{G}}_{A}$, its normalized MS score on $\phi$ is computed as:
@@ VI-B Experimental Results
Exp-1: Effectiveness. We evaluate the overall performance of the $\mathsf{GNN}$ explainers using QIs.
QI-1: IPF Scores. We report IPF scores (bigger is better) for all $\mathsf{GNN}$ explainers in Figure 3(a) with $\mathsf{GCNs}$ and kk=$10$. Additional explainability results with varying $k$ are given in the full version [47]. (1) In general, $\mathsf{ASX}$-$\mathsf{OP}$ outperforms a majority of competitors in multi-objective explainability evaluation metrics. For example, on $\mathsf{Cora}$, $\mathsf{ASX}$-$\mathsf{OP}$ outperforms $\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, and $\mathsf{SAME}$ in IPF scores by 1.34, 1.51, 2.81, 1.79, and 1.23 times, respectively. $\mathsf{DSX}$ and $\mathsf{ASX}$-$\mathsf{I}$ achieve comparable performance with $\mathsf{ASX}$-$\mathsf{OP}$. (2) We also observe that $\mathsf{GExp}$ and $\mathsf{PGExp}$ are sensitive as the datasets vary. For example, on $\mathsf{FacebookPage}$, both show a significant change in IPF scores. In contrast, $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ consistently achieve top IPF scores over all datasets. This verifies their robustness in generating high-quality explanations over different data sources.
@@
QI-3: nMS Scores. Figure 3(c) visualizes the normalized MS scores of the $\mathsf{GNN}$ explainers, on $\mathsf{Cora}$ dataset, where kk=$10$. (1) $\mathsf{ASX}$-$\mathsf{OP}$ reports a large range and contributes to explanations that are either close or the exact optimal explanation, for each of the three measures. $\mathsf{ASX}$-$\mathsf{I}$ and $\mathsf{DSX}$ have comparable performance, yet with larger ranges due to its properly diversified solution. (2) $\mathsf{DSX}$, $\mathsf{ASX}$-$\mathsf{OP}$, and $\mathsf{ASX}$-$\mathsf{I}$ contribute to the optimal $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$}, respectively. On the other hand, each individual explainer performs worse for all measures, with a large gap. For example, in $\mathsf{Cora}$, $\mathsf{MOExp}$ only achieves up to 3% of the best explanation ($\mathsf{DSX}$) over $\mathsf{fdl}$$^{-}$.
IPF scores over subsets of $\Phi$. Figure 3(d) presents a lattice graph where each node represents a subset of explainability measures ($\Phi$), and each edge corresponds to the removal of one measure $\phi$ from a top-down perspective. When $|\Phi|$=11, we observe that $\mathsf{conc}$} and $\mathsf{fdl}$$^{-}$ achieve the highest IPF scores, suggesting that they are the most discriminative measures in terms of dominance. In contrast, $\mathsf{acc}$} and $\mathsf{fdl}$$^{+}$ obtain relatively low IPF scores, indicating weaker dominance capability. For instance, the subset ${\mathsf{fdl}}^{+},{\mathsf{acc}}$} yields the lowest IPF score of $0.20$, whereas the subset ${\mathsf{fdl}}^{-},{\mathsf{conc}}$} achieves the highest score $0.94$, reflecting stronger dominance power compared to the other two measures.
@@
(1) $\mathsf{ASX}$-$\mathsf{OP}$, $\mathsf{ASX}$-$\mathsf{I}$, and $\mathsf{DSX}$ outperform all (learning-based) $\mathsf{GNN}$ explanations. $\mathsf{ASX}$-$\mathsf{OP}$ (resp. $\mathsf{DSX}$) on average outperforms $\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, and $\mathsf{SAME}$ by 2.05, 4.46, 9.73, 6.81, and 3.17 (resp. 1.62, 3.61, 7.88, 5.51, and 2.57) times, respectively. Moreover, $\mathsf{ASX}$-$\mathsf{OP}$ surpass $\mathsf{GExp}$, $\mathsf{PGExp}$, and $\mathsf{SAME}$ over larger graphs. $\mathsf{CF\textsuperscript{2}}$ and $\mathsf{MOExp}$ fail to generate explanations due to high memory cost and long waiting time. Learning costs remain their major bottleneck.
@@
(b) Dominance Coverage of Skyline Explanation v.s. Baseline Explanation ($\mathsf{AmazonComputer}$).
@@
(d) Dominance Coverage of Skyline Explanation v.s. Baseline Explanation ($\mathsf{FacebookPage}$).
Figure 5: Qualitative analysis of our skyline explanation and baseline explainers ($\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, $\mathsf{SAME}$).
Exp-3: Scalability. We report the impact of critical factors (i.e., number of explanatory subgraphs $k$, $\mathsf{GNN}$ classes, and number of $\mathsf{GNN}$ layers $L$) on the scalability of skyline explanation generation, using $\mathsf{Cora}$ dataset. Additional results about effectiveness by varying factors are in full version [47].
@@
Varying $\mathsf{GNN}$ classes and $L$. (1) Fixing kk=$10$, we report the time cost of $3$-layer $\mathsf{GNN}$ explainers for $\mathsf{GCN}$, $\mathsf{GAT}$, and $\mathsf{GIN}$ over $\mathsf{Cora}$. As shown in Figure 4(c), all our skyline methods take the least time to explain $\mathsf{GNN}$-based classification. This is consistent with our observation that the verification of $\mathsf{GNN}$ is the most efficient among all classes, indicating an overall small verification cost. (2) We fix kk=$10$ and report the time cost of $\mathsf{GCN}$ explainers, with the number of layers $L$ varied from 1 to 3 over $\mathsf{Cora}$. Figure 4(d) shows us that all our methods significantly outperform the competitors. The learning overheads of competitors remain their major bottleneck, while our algorithms, as post-hoc explainers without learning overhead, are more efficient. As expected, all our methods take a longer time to generate explanations for larger $L$, as more subgraphs need to be verified from larger induced $L$-hop neighbors.
@@
In Figure 5(b), we visualize the distribution of explanatory subgraphs for $v_{1}$ in Figure 5(a). Each point represents a subgraph with 3D coordinates given by $\mathsf{fdl}$$^{+}$, $\mathsf{fdl}$$^{-}$, and $\mathsf{conc}$} scores. The gray dots denote the verified explanatory subgraph in the interpretable space, while baseline explainers ($\mathsf{GExp}$, $\mathsf{PGExp}$, $\mathsf{CF\textsuperscript{2}}$, $\mathsf{MOExp}$, $\mathsf{SAME}$) are shown as diamonds. Our skyline explanations appear as red crosses, with their convex hull illustrating dominance coverage. We observe that skyline explanations span most of the interpretable space, offering comprehensive and diverse explanations, whereas baseline explanations cluster in smaller regions, missing diversity [18].
Skyline vs. Multi-objective Explanation (Linear Combination). Our second case compares $\mathsf{DSX}$ and $\mathsf{PGExp}$ (shown in Figure 5(c)). The latter generates explanations by optimizing a single objective that combines two explanation measures (factuality and conciseness). We observe that $\mathsf{PGExp}$ generates small factual subgraphs, such as $G_{\zeta}^{11}$ for a test node $v_{2}$, yet relatively less informative. $\mathsf{DSX}$ generates a skyline explanation $\{G_{\zeta}^{12},G_{\zeta}^{13}\}$ that is able to interpret $v_{2}$ as “Politicians” not only due to its connection with “Governments” entities but also highly related to “TV shows” and “Companies”. Meanwhile, Figure 5(d) shows that our skyline explanations span most of the interpretable space, providing diverse dominance coverage, whereas baseline explainers cluster their solutions within limited regions.